OpenStudy (anonymous):

evaluate the following without using a calculator. 7. arccot(sqrt3/3) 8. arcsec(2)

myininaya (myininaya):

So in number 7, there isn't a type-0? For number 8, let y=arcsec(2) So sec(y)=2 => cos(y)=1/2 When is cos(y)=1/2?

myininaya (myininaya):

Use unit circle.

OpenStudy (jdoe0001):

\(\bf cot^{-1}\left(\frac{\sqrt{3}}{3}\right)\implies \theta\implies cot(\theta)=\left(\frac{\sqrt{3}}{3}\right)\\ \quad \\ tan(\theta)=\left(\frac{3}{\sqrt{3}}\right)\implies tan(\theta)=\sqrt{3}\) notice the value is positive, and the constraints of the inverse cotangent function will put it in the 1st Quadrant

OpenStudy (anonymous):

thanks !

myininaya (myininaya):

Oh for number 7 you could make a right triangle and figure out and look at cosine and evaluate arccot(sqrt(3)/3) that way.